OpenStudy (anonymous):

Did you know that if you divide a number A of n digits (where the numerator is less than the denominator) by a string if n nines........

OpenStudy (anonymous):

That you get .AAAAAAAAAAAAAAAAA as an answer?

OpenStudy (kainui):

1/3=.333333 2/3=.666666 3/3=.999999

OpenStudy (anonymous):

63/99 = .6363636363636363636363636363636363 100/999 = .100100100100100100100100100100100 8172/9999=.817281728172817281728172817281728172

OpenStudy (anonymous):

http://mathforum.org/library/drmath/view/55740.html